# Algebraic Control: Complete Stable Inversion with Necessary and Sufficient Conditions

### v1

@@ I Introduction
Recent advancements in trajectory planning, autonomous vehicles, and robotics have increased interest in stable inversion to meet performance demands [1]. Learning-based control methods have further broadened research into inversion [2, 3], emphasizing the need to re-examine the conditions for stable inversion.
@@
Recent advancements include stable inversion for SISO affine systems [10], MIMO square systems in continuous time [11] and discrete time [12], and approximate stable inversion [13]. Specialized approaches, such as strong inversion for minimum-phase systems [14] and geometric methods for convolution-based inversion [15], have further expanded the field.
As the field evolves, the need for system classification and tailored solutions becomes evident. The primary objective of inversion-based approaches is to determine a bounded input that accurately reproduces the desired or given output. However, this goal poses challenges, particularly with non-minimum phase systems, which complicate the establishment of stable inputs under causality without infinite pre-actuation [9]. Furthermore, stable inversion can impact tracking performance and stability in the presence of uncertainties [16]. Merging learning-based control with stable inversion addresses these issues [10, 17].
Stable inversion methods have been applied across various domains, including iterative learning control [18], trajectory tracking for autonomous systems [19], and optimal channel equalization [20].
In this paper, we extend the concept of Zero Phase Error Tracking (ZPET) [8] to continuous-time systems. Our study focuses on determining the necessary and sufficient conditions for stable inversion in non-square, non-minimum phase, and uncertain systems. The main contributions are as follows:
  • Revisiting the stable inversion problem using a similar algebraic setting in Exact Model Matching [21] yet explicitly stating the output structures for necessary and sufficient conditions, which also hold for non-minimum phase systems. Compared to [14, 13, 11], the exact inversion conditions for non-square and even singular systems are also established.
  • Proposing an algebraic approximation method based on $\mathcal{H}_{\infty}$ theory when perfect tracking conditions are not met. This method allows for near-perfect tracking without the need for preview or pre-actuation approaches, which differentiates it from the methods used in [9, 15].
  • Ensuring robustness guarantees without the need for learning mechanisms, complementing the strategy in [10]. Furthermore, we establish upper bounds on achievable performance for non-minimum phase uncertain systems, extending the results presented in [13].
Our ultimate goal is to establish the necessary and sufficient conditions for stable inversion for non-minimum and uncertain systems in a causal manner, providing a unified framework to address the challenges of stable inversion.
The remainder of the paper is organized as follows: Section II provides the algebraic preliminaries and problem formulation. Section III presents the main results, including exact and almost necessary and sufficient conditions, as well as robustness. Section IV illustrates numerical examples to validate the theory. Section V concludes the paper.
@@ II-A Algebraic Preliminaries
Let $\mathbb{R}$ be a field of real numbers. Consider the set of all polynomials in $s$ with coefficients in $\mathbb{R}$. This set forms a commutative ring over $\mathbb{R}$ and is denoted by $\mathbb{R}[s]$. If this ring, say $\mathcal{R}$, has an identity element and no zero divisors, then $\mathcal{R}$ is an integral domain. Note that $\mathbb{R}[s]$ is also an Euclidian Domain. Moreover, the set of all such rational functions in $s$ over $\mathbb{R}$ forms a field, denoted by $\mathbb{R}(s)$ [22], and $\mathbb{R}[s]\subset\mathbb{R}(s)$ holds. The sets of $n_{y}\times n_{u}$ matrices with elements in $\mathbb{R}$, $\mathbb{R}[s]$, and $\mathbb{R}(s)$ are denoted by $\mathbb{R}^{n_{y}\times n_{u}}$, $\mathbb{R}[s]^{n_{y}\times n_{u}}$, and $\mathbb{R}(s)^{n_{y}\times n_{u}}$ respectively. Any $n_{y}\times n_{u}$ matrix, say $\grave{P}(s)\in\mathbb{R}[s]^{n_{y}\times n_{u}}$, over $\mathcal{R}$ can be factorized (see [23, Theorem 2.1]) as
$\grave{P}(s)$ $={\color[rgb]{0,0,0}U_{R}(s)S_{m}(s)V_{R}(s)}$
@@
Let $V$ be a vector space over the field $\mathbb{R}(s)$ with dimension $k$, consisting of $n$-tuples such that a basis of vector polynomials can always be found for $V$ [page 22, [24]]. A minimal basis of $V$ is defined as a $k\times n$ polynomial matrix $P_{m}$ [22]. Adapted Forney’s Theorem [22, Section 3(4.)] states that if $y=xP_{m}$ is a polynomial $n$-tuple, then $x$ must be a polynomial $k$-tuple.
@@
Some further notations throughout the paper are: (.$\|(.)(t)\|_{\infty}\triangleq\text{ess sup}_{t}|(.)(t)|$, any complex number can be expressed as (.)±$\Re(.)\pm j\Im(.)\in\mathbb{C}$, $\sigma$ represents singular values. Time domain square-integrable functions are denoted by $\mathcal{L}_{2}(-\infty,\infty)$. Its causal subset is given by $\mathcal{L}_{2}[0,\infty)$. For the frequency domain, including at $\infty$, $\mathcal{L}_{2}(j\mathbb{R})$ represents square-integrable functions on $j\mathbb{R}$, and $\mathcal{L}_{\infty}(j\mathbb{R})$ denotes bounded functions on $\Re(s)=0$. All these functions in Lebesgue spaces may be either matrix-valued or scalar. Then, $\mathcal{R}\mathcal{H}_{\infty}$ denotes the set of real rational $\mathcal{L}_{\infty}(j\mathbb{R})$ functions analytic in $\Re(s)>0$, $(.)^{\dagger}$ represents pseudo-inverse yielding unit matrix under multiplication, $(.)^{*}$ denotes complex conjugates transpose, and .,$\langle$.,$\rangle$ . , . ⟩ denotes the inner product. Now, among various definitions for multivariable zeros of a Transfer Function Matrix (TFM), $P(s)$, we adopt the definition given in [26, chapter 4.5.3] as
@@ II-B Problem Statement
where $A\in\mathbb{R}^{n\times n}$, $B\in\mathbb{R}^{n\times n_{u}}$, $C\in\mathbb{R}^{n_{y}\times n}$, $x,x_{0}\in\mathbb{X}\subseteq\mathbb{R}^{n},x(t)\in\mathcal{L}_{2}[0,\infty),y\in\mathbb{Y}\subset\mathbb{R}^{n_{y}},y(t)\in\mathcal{L}_{2}[0,\infty),u\in\mathbb{U}\subset\mathbb{R}^{n_{u}}$, and $u(t)\in\mathcal{L}_{2}[0,\infty)$. Moreover, the TFM representation of $\Xi({\color[rgb]{0,0,0}u},0)\triangleq C(sI-A)^{-1}B=P(s)$ and since all real rational strictly proper transfer matrices with no poles on the imaginary axis form $\mathcal{RL}_{2}(j\mathbb{R})\subset\mathcal{L}_{2}(j\mathbb{R})$ [27, page 48], $P(s)$ belongs to $\mathcal{L}_{2}(j\mathbb{R})\cap\mathbb{R}(s)^{n_{y}\times n_{u}}$. The system, $\Xi$, is; P1) a non-minimum phase s.t. $Z_{R}\neq\emptyset$, P2) either $n_{y}=n_{u}$ (square) or $n_{y}\neq n_{u}$ (nonsquare), and casual ($t>0$).
@@
Under Assumption 1, let $y_{1}(t)$ be an observed output trajectory of the system $\Xi$ in (6) resulting from some unknown initial condition $\hat{x}_{0}$ and control input $\hat{u}$.
Now, given trajectory $y_{1}(t)$ and initial condition $x_{0}$, the left inverse, $\Xi_{i}^{\ell}:\mathbb{Y}\times\mathbb{X}\to\mathbb{U}$, is defined such that
$u_{inv}(t)=\Xi_{i}^{\ell}(y_{1}(t),x_{0}),\quad\|u_{inv}(t)\|_{\infty}<\infty.$ (7)
The error is
@@
The inverse, $\Xi_{i}^{\ell}$, is classified based on $e(t)$ as follows:
@@
For approximate inversion, among all possible approximations, our goal is to find $\Xi_{i}^{\ell}$ that minimizes the error $e(t)$ with $\|u_{inv}(t)\|_{\infty}<\infty$ and satisfying
@@
The terminology in Problem 1, adopted from [28], allows $y_{1}(t)$ to represent desired trajectories, as in [12]. Assumption 1 and $\|e(t)\|_{\infty}<\infty$ ensure Lyapunov stability for the error dynamics in (8). The objective of minimizing the error aligns with [13], where $\varphi$ in (9) accounts for bounded yet inevitable errors, including steady-state errors due to system limitations.
To solve Problem 1.a algebraically, given $y_{1}(t)$, define $y_{ic}(t)\triangleq\Xi(0,x_{0})$, the trajectory caused by the known initial conditions. To isolate $y_{ic}(t)$ in $y_{1}(t)$, we take the Laplace transform (see Paley-Wiener Th., [29, p.104] for the existence) yielding
[…]
@@
From this point on, similar to the [21, Theorem 8.5.2], Problem 1.a reduces to algebraically solving the rational matrix equation.
Remark 2.
The algebraic solution of (11) is not restricted solely to the framework of (10). It can also be expressed through describing functions [30, Sec III.C]. Furthermore, the recent work in [31] presents the Koopman resolvent, a Laplace-domain representation for specific nonlinear dynamics. This development emphasizes the broader potential for algebraic solutions in strongly nonlinear systems.
@@ III-A Left Inverse
$\qquad+(I_{n_{u}}-\hat{V}_{R1}{V}_{R1})\bar{\kappa}\Biggr{)}\grave{Y}(s)=\Xi_{i}^{\ell}(s)\grave{Y}(s)$ (14)
@@ III-B Algebraic Necessary and Sufficient Conditions
Consider $\Xi$ in (6) with P1-P2 and $\Xi^{\ell}_{i}$ in (III-A). Then sufficient and necessary condition for Problem 1.a. is
@@
$U(s)$ $=\hat{V}_{R1}\Lambda^{-1}\hat{U}_{R1}\grave{Y}(s)$
$\Lambda\hat{V}^{\ell}_{R1}U(s)$ $=\hat{U}_{R1}\grave{Y}(s)$
[…]
where there always exist $\hat{V}^{\ell}_{R1}$ satisfying $\hat{V}^{\ell}_{R1}\hat{V}_{R1}=I_{r}$. Now assume that either $\grave{P}(s)$ or $U_{R1}\Lambda\hat{V}^{\ell}_{R1}$ is not a minimal basis. The minimal basis $\grave{P}_{m}(s)$ of an $n_{y}$-dimensional vector space of $n_{u}$-tuples over $\mathbb{R}[s]$ for (13) is given by $\grave{P}_{m}(s)=L_{T}(s)\grave{P}(s)$ for condition C1, and $\grave{P}_{m}(s)=L_{T}(s)U_{R1}\Lambda\hat{V}^{\ell}_{R1}$ for condition C2. Here, the transformation $L_{T}(s)\in\mathbb{R}(s)^{n_{y}\times n_{y}}$, which yields the minimal basis, exists with $\det(L_{T})\neq 0$ (see [22, Section 2]). Then, left multiplying $L_{T}(s)$ to (13) or (16) results in
@@
where $U_{N}(s)\in\mathbb{R}[s]^{n_{u}}$ and scalar $U_{D}(s)\in\mathbb{R}{\color[rgb]{0,0,0}[s]}$. Now define a new scalar $z^{+}(s)$ which contains all RHP zeros of $\grave{P}(s)$ as
$z^{+}(s)=\prod_{z\in Z_{R}{\color[rgb]{0,0,0}\bigcap\mathbb{R}}}(s-z)^{m_{z}}\prod_{z\in{\color[rgb]{0,0,0}Z_{R}\bigcap(\mathbb{C}\setminus\mathbb{R})}}(s-z)^{m_{z}}(s-{z}^{*})^{m_{{z}^{*}}}$ (22)
where $z^{+}(s)\in\mathbb{R}[s]\subset\mathbb{R}(s)$, $m_{z}$ and $m_{z^{\ast}}$ denote the multiplicities of $z$ and ${z}^{*}$ respectively. Thus, there exist non-zero vector $y^{\ast}_{z}\ (\text{or }u_{z})$ such that $y^{\ast}_{z}\grave{P}(z)=0$ (or $\grave{P}(z)u_{z}=0.y_{z}$) $\forall z\in Z_{R}$.
@@
While the algebraic equation is inspired by Exact Model Matching for invertibility, we explicitly define necessary and sufficient conditions for the system’s output, applicable even to non-minimum phase systems by leveraging polynomial bases. Specifically, the condition $Y(s)\in\operatorname*{Im}_{\mathbb{R}(s)}(P)$ defines the set of outputs reachable through the input channel, while $Y(z)=0,\ \forall z\in Z_{R}$ redefines the output on $Z_{R}$. Relaxing either condition results in losing the property “$=0$” over the $e(t)$.
Notably, for certain specific conditions, such as minimum phase systems, the findings can be applied more directly.
@@
Assume $P(s)$ is square and full rank so that $U_{R}=U_{R1},V_{R}=V_{R1},(I_{n_{u}}-\hat{V}_{R1}{V}_{R1})=0$, and $Z_{R}=\emptyset$, then (III-A) becomes
$U(s)={\hat{V}_{R1}\Lambda^{-1}(s)\hat{U}_{R1}}Y(s).$ (25)
@@
Assume $P(s)$ is non-square and full rank, so that $V_{R}=V_{R1},(I_{n_{u}}-\hat{V}_{R1}{V}_{R1})=0$, and $Z_{R}=\emptyset$, then (III-A) becomes
$U(s)$ $=\begin{bmatrix}\hat{V}_{R1}\Lambda^{-1}(s)&\textbf{0}_{n_{u}\times(n_{y}-n_{u})}\end{bmatrix}\begin{bmatrix}\hat{U}_{R1}\\ \hat{U}_{R2}\end{bmatrix}Y(s)$ (26)
$=(P^{*}(s)P(s))^{-1}P^{*}(s)Y(s)=P^{\dagger}(s)Y(s).$ (27)

### v2

@@ I Introduction
Recent advancements in autonomous systems and robotics have increased interest in stable inversion to meet performance demands. Learning-based control methods have further broadened research into inversion [1], emphasizing the need to re-examine the conditions for stable inversion.
@@
Recent advancements include stable inversion for SISO affine systems [8], square systems in continuous time [9] and discrete time [10], and approximate inversion [11]. Additionally, strong inversion for handling initial states in multivariable systems [12] and geometric methods for convolution-based inversion [13] have further expanded the field.
As the field evolves, the need for system classification and corresponding solutions becomes evident. The primary objective of inversion-based approaches is to determine a bounded input that accurately reproduces the desired or given output. However, this goal poses challenges, particularly with non-minimum phase systems, which complicate the establishment of stable inputs under causality without infinite pre-actuation [7]. Furthermore, stable inversion can impact tracking performance and stability in the presence of uncertainties [14]. Merging learning-based control with stable inversion partially addresses these issues [8, 15].
Despite the theoretical challenges, the practical side is equally important. Stable inversion methods have been applied across various domains, including iterative learning control [16], trajectory tracking for autonomous systems [17], and optimal channel equalization [18].
In this paper, we extend the concept of Zero Phase Error Tracking (ZPET) [6] to continuous-time systems. In this setting, we may apply ZPET for compensating the phase shifts introduced by unstable zeros, but ZPET compensation alone does not make sense due to the high-pass nature of these zeros. However, if the overall compensator is designed to suppress these high-pass effects, ZPET remains effective at low frequencies, enabling near-perfect tracking of low-frequency sinusoidal references. The main contributions are:
1) We revisit the stable inversion problem using a similar algebraic setting of Model Matching [19], yet specifying output structures leading to necessary and sufficient conditions, even for non-minimum phase, non-square, and singular systems.
2) We propose an algebraic approximation that repurposes certain functions in $\mathcal{H}_{\infty}$ theory to achieve near-perfect tracking when perfect tracking conditions are not met, eliminating the need for preview or pre-actuation.
3) Our approach guarantees robustness without learning mechanisms. When limitations or uncertainties make the output set only partially reachable, we prove that the tracking error converges to an inevitable yet a bounded residual.
Our goal is to develop a unified, causal framework for stable inversion in non-minimum phase, multivariable, and uncertain systems, addressing the inherent challenges.
We organize this paper as follow: Section II provides the preliminaries. Section III presents the main results. Section IV illustrates numerical examples. Section V concludes the paper.
@@ II-A Algebraic Preliminaries
Let $\mathbb{R}$ be a field of real numbers. Consider the set of all polynomials in $s$ with coefficients in $\mathbb{R}$. This set forms a commutative ring over $\mathbb{R}$ and is denoted by $\mathbb{R}[s]$. If this ring, say $\mathcal{R}$, has an identity element and no zero divisors, then $\mathcal{R}$ is an integral domain. Note that $\mathbb{R}[s]$ is also a Euclidean Domain. Moreover, the set of all such rational functions in $s$ over $\mathbb{R}$ forms a field, denoted by $\mathbb{R}(s)$ [20], and $\mathbb{R}[s]\subset\mathbb{R}(s)$ holds. The sets of $n_{y}\times n_{u}$ matrices with elements in $\mathbb{R}$, $\mathbb{R}[s]$, and $\mathbb{R}(s)$ are denoted by $\mathbb{R}^{n_{y}\times n_{u}}$, $\mathbb{R}[s]^{n_{y}\times n_{u}}$, and $\mathbb{R}(s)^{n_{y}\times n_{u}}$ respectively. Any $n_{y}\times n_{u}$ matrix, say $\grave{P}(s)\in\mathbb{R}[s]^{n_{y}\times n_{u}}$, over $\mathcal{R}$ can be factorized (see [21, Theorem 2.1]) as
$\grave{P}(s)$ $=U_{R}(s)S_{m}(s)V_{R}(s)$
@@
Let $V$ be a vector space over the field $\mathbb{R}(s)$ with dimension $k$, consisting of $n$-tuples such that a basis of vector polynomials can always be found for $V$ [page 22, [22]]. A minimal basis of $V$ is defined as a $k\times n$ polynomial matrix $P_{m}$ [20]. Adapted Forney’s Theorem [20, Section 3(4.)] states that if $y=xP_{m}$ is a polynomial $n$-tuple, then $x$ must be a polynomial $k$-tuple.
@@
Some further notations throughout the paper are: (.$\|(.)(t)\|_{\infty}\triangleq\text{ess sup}_{t}|(.)(t)|$, any complex number can be expressed as (.)±$\Re(.)\pm j\Im(.)\in\mathbb{C}$, $\sigma$ represents singular values. Time domain square-integrable functions are denoted by $\mathcal{L}_{2}(-\infty,\infty)$. Its causal subset is given by $\mathcal{L}_{2}[0,\infty)$. For the frequency domain, including at $\infty$, $\mathcal{L}_{2}(j\mathbb{R})$ represents square-integrable functions on $j\mathbb{R}$, and $\mathcal{L}_{\infty}(j\mathbb{R})$ denotes bounded functions on $\Re(s)=0$. All these functions in Lebesgue spaces may be either matrix-valued or scalar. Then, $\mathcal{R}\mathcal{H}_{\infty}$ denotes the set of real rational $\mathcal{L}_{\infty}(j\mathbb{R})$ functions analytic in $\Re(s)>0$, $(.)^{\dagger}$ represents pseudo-inverse yielding unit matrix under multiplication, $(.)^{*}$ denotes complex conjugates transpose, and .,$\langle$.,$\rangle$ . , . ⟩ denotes the inner product. Now, among various definitions for multivariable zeros of a Transfer Function Matrix (TFM), $P(s)$, we adopt the definition given in [24, chapter 4.5.3] as
@@ II-B Problem Statement
where $A\in\mathbb{R}^{n\times n}$, $B\in\mathbb{R}^{n\times n_{u}}$, $C\in\mathbb{R}^{n_{y}\times n}$, $x,x_{0}\in\mathbb{X}\subseteq\mathbb{R}^{n},x(t)\in\mathcal{L}_{2}[0,\infty),y\in\mathbb{Y}\subset\mathbb{R}^{n_{y}},y(t)\in\mathcal{L}_{2}[0,\infty),u\in\mathbb{U}\subset\mathbb{R}^{n_{u}}$, and $u(t)\in\mathcal{L}_{2}[0,\infty)$. Moreover, the TFM representation of $\Xi(u,0)\triangleq C(sI-A)^{-1}B=P(s)$ and since all real rational strictly proper transfer matrices with no poles on the imaginary axis form $\mathcal{RL}_{2}(j\mathbb{R})\subset\mathcal{L}_{2}(j\mathbb{R})$ [25, page 48], $P(s)$ belongs to $\mathcal{L}_{2}(j\mathbb{R})\cap\mathbb{R}(s)^{n_{y}\times n_{u}}$. The system, $\Xi$, is; P1) a non-minimum phase s.t. $Z_{R}\neq\emptyset$, P2) either $n_{y}=n_{u}$ (square) or $n_{y}\neq n_{u}$ (nonsquare), and casual ($t>0$).
@@
Under Assumption 1, let $\Xi$ denotes the forward system given in (6), and let $y_{1}(t)$ be an observed output trajectory of $\Xi$. We define the right inverse $\Xi_{i}^{\Gamma}:\mathbb{Y}\times\mathbb{X}\to\mathbb{U}$, which, given $y_{1}(t)$ and a prescribed inverse system’s initial states, $\bar{x}_{0}\in\mathbb{X}$, produces the control input
$u_{inv}(t)=\Xi_{i}^{\Gamma}(y_{1}(t),\bar{x}_{0}),\quad\|u_{inv}(t)\|_{\infty}<\infty$ (7)
Let $x_{0}$ be the initial states of the forward system. The error is then defined by
@@
The inverse, $\Xi_{i}^{\Gamma}$, is classified based on $e(t)$ as follows:
@@
For approximate inversion, among all possible approximations, our goal is to find $\Xi_{i}^{\Gamma}$ that minimizes $e(t)$ and satisfies
@@
The terminology in Problem 1, adopted from [26], allows $y_{1}(t)$ to represent desired trajectories, as in [10]. Assumption 1 and $\|e(t)\|_{\infty}<\infty$ ensure Lyapunov stability for the error dynamics in (8). While our objective of minimizing $e(t)$ is in line with [11], we further introduce a residual term $\varphi$ in (9) capturing the inevitable error due to system limitations.
Remark 2.
As is typical in exact stable inversion, we initially assume that $(A,B,C)$ in (6) and the initial states $x_{0}$ and $\bar{x}_{0}$ are known. In Section III-D, this assumption is relaxed to accommodate model uncertainty, and in Sections III-C-III-D, the requirement on $x_{0}$ and $\bar{x}_{0}$ is fully removed, thus allowing unknown initial states but resulting in only an approximate solution. For further details on initial states, see [12].
To solve Problem 1.a algebraically, define $y_{ic}(t)\triangleq\Xi(0,x_{0})$, $y\triangleq y_{1}-y_{ic}$, $u_{inv}^{(0)}(t)\triangleq\Xi_{i}^{\Gamma}(0,\bar{x}_{0})$, and $u(t)\triangleq\Xi_{i}^{\Gamma}(y(t),0)$ to represent the trajectories and control inputs for known or zero initial states. Then, Remark 2 let us re-define (7) as $u_{inv}(t)\triangleq\Xi_{i}^{\Gamma}(y_{1}(t),\bar{x}_{0})-u_{inv}^{(0)}(t)$. Applying the Laplace transform (see Paley–Wiener Th. [27, p. 104] for existence) yields
$\mathscr{L}\{y_{1}(t)-{y}_{ic}(t)\}={C}(s{I}-{A})^{-1}{B}U(s)=Y(s)$ (10)
@@
From this point on, similar to the [19, Theorem 8.5.2], Problem 1.a reduces to algebraically solving the rational matrix equation where $\|u(t)\|_{\infty}<\infty\iff\|u_{inv}(t)\|_{\infty}<\infty$.
@@ III-A Right Inverse
$\qquad+(I_{n_{u}}-\hat{V}_{R1}{V}_{R1})\bar{\kappa}\Biggr{)}\grave{Y}(s)=\Xi_{i}^{\Gamma}(s)\grave{Y}(s)$ (14)
@@
Remark 3.
From an algebraic standpoint, although obtaining $U(s)$ involves multiplying $\grave{P}(s)$ by its left inverse, following [5, 26], we denote $\Xi_{i}^{\Gamma}$ as a “right inverse”.
@@ III-B Algebraic Necessary and Sufficient Conditions
Consider $\Xi$ in (6) with P1-P2 and $\Xi^{\Gamma}_{i}$ in (III-A). Then sufficient and necessary condition for Problem 1.a. is
@@
$U(s)=\hat{V}_{R1}\Lambda^{-1}\hat{U}_{R1}\grave{Y}(s)\implies\Lambda\hat{V}^{\ell}_{R1}U(s)=\hat{U}_{R1}\grave{Y}(s)$
$\implies U_{R1}\Lambda\hat{V}^{\ell}_{R1}U(s)=U_{R1}\hat{U}_{R1}\grave{Y}(s)=\grave{Y}(s)$ (16)
where there always exist $\hat{V}^{\ell}_{R1}$ satisfying $\hat{V}^{\ell}_{R1}\hat{V}_{R1}=I_{r}$. Now assume that either $\grave{P}(s)$ or $U_{R1}\Lambda\hat{V}^{\Gamma}_{R1}$ is not a minimal basis. The minimal basis $\grave{P}_{m}(s)$ of an $n_{y}$-dimensional vector space of $n_{u}$-tuples over $\mathbb{R}[s]$ for (13) is given by $\grave{P}_{m}(s)=L_{T}(s)\grave{P}(s)$ for condition C1, and $\grave{P}_{m}(s)=L_{T}(s)U_{R1}\Lambda\hat{V}^{\Gamma}_{R1}$ for condition C2. Here, the transformation $L_{T}(s)\in\mathbb{R}(s)^{n_{y}\times n_{y}}$, which yields the minimal basis, exists with $\det(L_{T})\neq 0$ (see [20, Section 2]). Then, left multiplying $L_{T}(s)$ to (13) or (16) results in
@@
where $U_{N}(s)\in\mathbb{R}[s]^{n_{u}}$ and scalar $U_{D}(s)\in\mathbb{R}[s]$. Now define a new scalar $z^{+}(s)$ which contains all RHP zeros of $\grave{P}(s)$ as
$z^{+}(s)=\prod_{z\in Z_{R}\bigcap\mathbb{R}}(s-z)^{m_{z}}\prod_{z\in Z_{R}\bigcap(\mathbb{C}\setminus\mathbb{R})}(s-z)^{m_{z}}(s-{z}^{*})^{m_{{z}^{*}}}$
where $z^{+}(s)\in\mathbb{R}[s]$, $m_{z}$ and $m_{z^{\ast}}$ denote the multiplicities of $z$ and ${z}^{*}$ respectively. Thus, there exist non-zero vector $y^{\ast}_{z}\ (\text{or }u_{z})$ s.t. $y^{\ast}_{z}\grave{P}(z)=0$ (or $\grave{P}(z)u_{z}=0.y_{z}$) $\forall z\in Z_{R}$.
@@
While the algebraic equation is inspired by Exact Model Matching, we explicitly define necessary and sufficient conditions on the system’s output, applicable even to non-minimum phase and singular systems. Specifically, the condition $Y(s)\in\operatorname*{Im}_{\mathbb{R}(s)}(P)$ defines the set of reachable outputs, while $Y(z)=0$ for all $z\in Z_{R}$ defines the set of outputs for which a corresponding stable input exists. Relaxing either condition results in losing the property ‘$=0$” over the $e(t)$.
Remark 5.
For the counterpart involving left inverses in contexts such as fault detection, see [12, 26], which provide necessary and sufficient conditions for invertibility under the assumptions of full-rank $P(s)$, $D$ and with $Z_{R}=\emptyset$.
Assuming $Z_{R}=\emptyset$ yields the following corollaries:
@@
Assume $P(s)$ is square and full rank so that $U_{R}=U_{R1},V_{R}=V_{R1},(I_{n_{u}}-\hat{V}_{R1}{V}_{R1})=0$, and $Z_{R}=\emptyset$, then (III-A) becomes $U(s)={\hat{V}_{R1}\Lambda^{-1}(s)\hat{U}_{R1}}Y(s)$.
@@
Assume $P(s)$ is non-square and full rank, so that $V_{R}=V_{R1},(I_{n_{u}}-\hat{V}_{R1}{V}_{R1})=0$, and $Z_{R}=\emptyset$, then (III-A) becomes $U(s)=\begin{bmatrix}\hat{V}_{R1}\Lambda^{-1}(s)&\textbf{0}_{n_{u}\times(n_{y}-n_{u})}\end{bmatrix}\begin{bmatrix}\hat{U}_{R1}\\ \hat{U}_{R2}\end{bmatrix}Y(s)=(P^{*}(s)P(s))^{-1}P^{*}(s)Y(s)=P^{\dagger}(s)Y(s)$.